OpenStudy (anonymous):

Solve the Inequalities: -k/3 + 8 < 7

OpenStudy (anonymous):

This is the real one: \[-k/3+8\le7\]

OpenStudy (callisto):

\[ -\frac{k}{3}+8\le7\]\[ -\frac{k}{3}+8-8\le7-8\]\[ -\frac{k}{3}\le-1\]\[ -\frac{k}{3} \times(-3)\ge-1\times(-3)\]\[ k\ge3\]

OpenStudy (anonymous):

what does that mean?

OpenStudy (callisto):

What do you mean? You can't see the latex?

OpenStudy (callisto):

Or you don't understand ?

OpenStudy (anonymous):

the first one....i think

OpenStudy (callisto):

(-k/3) + 8 ≦7 (-k/3) + 8 -8≦ 7-8 (-k/3) ≦ -1 (-k/3) x (-3) ≧ -1 x (-3) k ≧ 3 Got it?

OpenStudy (anonymous):

kind of

OpenStudy (callisto):

Hmm.. Which part you don't understand?

OpenStudy (anonymous):

i think i got it im just looking at it harder

OpenStudy (anonymous):

i get it up to step 5

OpenStudy (callisto):

lol Step 5 is the last step...

OpenStudy (anonymous):

i ment 4 sorry

OpenStudy (callisto):

You don't understand the change of sign?

OpenStudy (anonymous):

nevermind i got it, thanks so much!!

OpenStudy (callisto):

Glad to know that you understand. You're welcome. Ask if you have any questions :)

OpenStudy (anonymous):

well do